You have a four-year financial instrument that will pay $106 in year 1, $215 in year 2, $300 in year 3, and $428 in year 4. If the appropriate discount rate is 17% p.a., find the duration of this instrument.
Answer #1


Period

Cash Flow from Bond

Discounting factor = 1/(1+R)^N

PV of the cash flows = Cash flow x Df

Weighted cash flow = Period x Cash flow

Present value of weighted cash flow = Weighted Cash flow x Df

N

CF

Df = 1/(1+17%)^N

PV = CF x Df

WCF = CF x N

WPV = WCF x Df

1

106.00

0.8547

90.5983

106.0000

90.5983

2

215.00

0.7305

157.0604

430.0000

314.1208

3

300.00

0.6244

187.3112

900.0000

561.9335

4

428.00

0.5337

228.4022

1712.0000

913.6089

Total = P = Price =

663.3721

Total = Weighted Price = WP

1880.2615

Duration or Macaulay Duration = WP/P

Duration or Macaulay Duration = 1880.2615/663.3721

Duration or Macaulay Duration = 2.834399427

Duration or Macaulay Duration = 2.834 years